Please clear handwriting and highlight answers thank you!!
9. (20 pts) A fishing line manufacturer claims that his new X-25 series has an average tensile strength (H) of 25 psi following a normal distribution. To verify his claim, a sample of 25 specimens was tested and yielded the following results: 32, 19, 27, 24, 28, 29, 38, 30, 25, 33, 18, 25, 37,28, 20, 32, 23,21, 25, 19, 23,27, 32, 26,30 this fishing line not meeting the specifications? Based on the 90% confidence interval for μ is the manufacturers claim valid? How large a sample is needed a. Is the sample data following a normal distribution? If the specification is 25 + 5 psi, whats the percentage of b. if we wish to be 90% confident that our sample mean will be within 1 psi of the true mean?
Answer #1

(a)

Following is the box plot of the data:

Descriptive statistics count mean sample standard deviation sample variance minimum maximum range Data 25 26.84 5.41 29.31 38 20 1st quartile median 3rd quartile interquartile range mode 23.00 27.00 30.00 7.00 32.00 low extremes low outliers high outliers high extremes 0 0 0 0 BoxPlot 15 20 25 30 35 40 Data

Following is the normal quantile plot of the data:

Normal Quantile Plot 2.5 1.5 0.5 0 -0.5 -1 1.5 -2 -2.5 10 1520 245 30 3540 Average Tensile strength

The box plot is symmetric and normal quantile plot shows approximate straight line pattern so we can assume that data is normally distributed.

The interval corresponding to 25 +/- 5 psi is (20, 30). Out of 25 data values, 16 lies in this interval and 25-16 = 9 does not lie in this interval. So the percentage of finishing time not meeting the specifications is

(9/25) *100% = 36%

-----------------------------------------

(b)

The 90% confidence interval.

Here we have

n-25,26.84, s5.41 rl

Since there 16 data values in the sample so degree of freedom is df=25-1=24 and critical value of t for 90% confidence interval, using excel function "=TINV(0.10,24)" is 1.711.

The required confidence interval is

5.41 ± t 26.84 ± 1.711 . = 26.84 ± 1.85 = (24.99, 28.69) critical Vn V25 っ

Therefore, a 90% confidence interval for the mean is (24.99, 28.69).

Since 25 lies in the above interval so it seems that the manufacturer's claim is valid.

-------------------

Assuming sample standard deviation is a good estimate of population standard deviation we have

sigma=5.41

Critical value of z for 90% confidence interval , α = 0.10 , is za/2 = 1.645 . The margin of error is E = 1 so required sample size will be

rl E」 =11.645-24112

So required sample size for 90% confidence interval is 80.